OpenStudy (anonymous):

4 divided by radical 6 minus radical 2

OpenStudy (unklerhaukus):

\[4\div\sqrt 6-\sqrt 2\]?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

would you mind drawing it out for me?

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

\[4÷\sqrt 6−\sqrt 2=\frac4{\sqrt 6}-\sqrt 2\]\[\qquad=\frac4{\sqrt 6}\times1-\sqrt 2\]\[\qquad=\frac4{\sqrt 6}\times\frac{\sqrt 6}{\sqrt 6}-\sqrt 2\]\[\qquad=\]

OpenStudy (anonymous):

what does it equal?